# Multi-Modal Recommendation Unlearning for Legal, Licensing, and Modality Constraints

### v2

@@ Abstract
This paper introduces MMRecUn, the first approach known to us for unlearning in MMRS and unlearning item data. Given a trained RS model, MMRecUn employs a novel Reverse Bayesian Personalized Ranking (BPR) objective to enable the model to forget marked data. The reverse BPR attenuates the impact of user-item interactions within the forget set, while the forward BPR reinforces the significance of user-item interactions within the retain set. Our experiments demonstrate that MMRecUn outperforms baseline methods across various unlearning requests when evaluated on benchmark MMRS datasets. MMRecUn achieves recall performance improvements of up to $\mathbf{49.85\%}$ compared to baseline methods and is up to $\mathbf{1.3}\times$ faster than the Gold model, which is trained on retain set from scratch. MMRecUn offers significant advantages, including superiority in removing target interactions, preserving retained interactions, and zero overhead costs compared to previous methods. The code will be released after review.
@@ 1 Introduction
Background and Related Work. Given these benefits, one might ask: Can unlearning methods for uni-modal systems be adapted for the unique challenges of multi-modal systems? Previous unlearning methods (Nguyen et al. 2022) face the following challenges: 1 MMRS integrate diverse user-item data, including images, text, and behavior, into a unified convolutional graph. This is incompatible with the matrices, latent factors, feature representations, and temporal sequences used in matrix factorization (Liu et al. 2023a; Xu et al. 2022; Liu et al. 2022; Zhang et al. 2023a), collaborative filtering (Li et al. 2024; Schelter, Ariannezhad, and de Rijke 2023), and sequential RS (Ye and Lu 2023), respectively. 2 For methods that do use graph-based representations, integrating diverse modalities is challenging. Because structure and feature embeddings are tightly integrated, unlearning in one modality has cascading effects on the other (Cheng and Amiri 2023). 3 Approximate unlearning methods (You et al. 2024; Zhang et al. 2023b; Li et al. 2023b) involve expensive operations, such as inverting Hessian matrix or the Fisher Information Matrix. These computations are costly and impractical for MMRS having a large number of feature dimensions. 4 Exact unlearning methods (Chen et al. 2022a; Li et al. 2023a) adapt SISA (Bourtoule et al. 2021) to split the dataset into shards. This disrupts the graph structure and degrades performance within each shard due to limited data and poor data heterogeneity (Koch and Soll 2023) by $10-30\%$ (You et al. 2024). The aggregation step, to preserve data structure and aggregate performance across shards, introduces significant overhead costs, impacting both training and inference, increasing proportionally with the number of shards (Ramezani et al. 2021). In worst case, when data that needs to be forgotten come from multiple shards, efficiency falls to the level of retraining from scratch. 5 While unlearning might appear efficient by retraining only specific shards, the process of setting up and aggregating these shards introduces additional, unaccounted overhead. If the data is initially partitioned by user dimensions, it cannot be re-split for item unlearning. Consequently, simultaneous unlearning of both user and item data becomes impossible.
These constraints underscore the need for methods specifically designed for unlearning in MMRS and for unlearning item data. The supplementary material (Appendix  A.1) provides further insights into the challenges and other previous works  (Cheng and Amiri 2023; Yuan et al. 2023; Li et al. 2023c; Chen et al. 2024; Ganhör et al. 2022; Xin et al. 2024; Wang et al. 2024; Sinha, Mandal, and Kankanhalli 2023; Tarun et al. 2023b; Chundawat et al. 2023b; Tarun et al. 2023a; Chundawat et al. 2023a).
@@ 3 Proposed MMRecUn Method
Traditionally, there are two stages of model $M(\cdot,\varphi)$. In training, the model encodes collaborative signals inherent in the user-item interaction matrix $\mathcal{Y}$. The optimization process minimizes the BPR loss (Rendle et al. 2012):
@@
Time Complexity for unlearning LightGCN with BPR and auxiliary contrastive loss using MMRecUn is $T=\mathcal{O}\left(\mathcal{M}\cdot d\cdot|\mathcal{E}|\cdot d_{m}+|\mathcal{U}|\cdot|\mathcal{I}|\cdot K+d^{2}+|\mathcal{M}|\cdot d\right)$, where $\mathcal{M}$ is the number of modalities, $d$ is the embedding dimension, $|\mathcal{E}|$ is the number of edges, $d_{m}$ is the feature dimension for modality $m$, $|\mathcal{U}|$ is the number of users, $|\mathcal{I}|$ is the number of items, and $K$ is the number of negative samples per user-item pair. See Appendix A.5 for derivation.
@@ 4.1 Experimental Setup
Datasets and baselines. We select Mgcn (Yu et al. 2023) as our base model for several reasons: 1 It is the state-of-the-art MMRS with BPR. 2 Mgcn represents BPR-based models well, making improvements likely to generalize. 3 Unlike traditional methods like Collaborative Filtering and Matrix Factorization, it handles multi-modal features through graph convolutional representations. 4 Other loss functions for uni-modal systems don’t address multi-modal complexities, and those that do, perform weaker than Mgcn. So, focusing on Mgcn ensures our approach is relevant.
We use three distinct categories within the Amazon dataset (Hou et al. 2024) that are used to benchmark MMRS (see Appendix A.6). The model trained on the retain set (named as the Gold model) from scratch is used as baseline. We prepare another baseline by re-purposing the image classification unlearning AmUn (Graves, Nagisetty, and Ganesh 2021) for MMRS.
@@
Evaluation metrics To evaluate a given model $M(\cdot,\varphi)$, we need 1 model predictions: the ranked list of user-item pairs; 2 the ground truth: user-item interactions $\mathcal{Y}$; and, 3 $K$: the number of the top recommendations to consider. To capture a variety of aspects of performance, we use 1 predictive metrics, that reflect the “correctness”, how well $M(\cdot,\varphi)$ finds relevant items such as Recall at K and Precision at K; and 2 ranking metrics, that reflect the ranking quality, how well $M(\cdot,\varphi)$ can sort the items from more relevant to less relevant. NDCG considers both the relevance and the position of items in the ranked list. MAP measures the average Precision across different Recall levels for a ranked list.
Settings and Hyper-parameters All the experiments are performed on $4$x NVIDIA RTX2080 ($32$GB). We use the identical settings as in the case of Mgcn for training the original model $M(\cdot,\varphi)$. The data interaction history of each user is split $8:1:1$ for training, validation and testing. The training set is used to create retain set and forget set. We initialize the embedding with Xavier initialization of dimension $64$, set the regularization coefficient to $\lambda_{\varphi}=10^{-4}$, and the batch size to $B=2048$. For the self-supervised task, we set the temperature to $0.2$. For convergence, the early stopping and total epochs are fixed at $20$ and $1000$, respectively. We use Recall@20 on the validation data as the training-stopping indicator following (Zhang et al. 2022). We conducted experiments five times using different seed values and report the standard deviation values for item unlearning in Table 3. For the other types of unlearning, the standard deviation values were found to be negligible and are therefore omitted.
@@
User-Centric Item-Centric
$\text{Recall}_{u}@K=\frac{|\text{Relevant}_{u}\cap\text{Retrieved}_{u}|}{|\text{Relevant}_{u}|}$ $\text{Recall}_{i}@K=\frac{\sum_{j=1}^{K}\text{rel}_{j,i}}{|\text{Users}_{i}|}$
$\text{Recall}_{\text{user}}@K=\frac{1}{|\mathcal{U}|}\sum_{u\in\mathcal{U}}\text{Recall}_{u}@K$ $\text{Recall}_{\text{item}}@K=\frac{1}{|\mathcal{I}|}\sum_{i\in\mathcal{I}}\text{Recall}_{i}@K$
$\text{Prec}_{u}@K=\frac{|Rel_{u}\cap Rec_{u}|}{|Rec_{u}|}$ $\text{Prec}_{i}@K=\frac{\sum_{j=1}^{K}\text{rel}_{j,i}}{K}$
$\text{Precision}_{\text{user}}@K=\frac{1}{|\mathcal{U}|}\sum_{u\in\mathcal{U}}\text{Prec}_{u}@K$ $\text{Precision}_{\text{item}}@K=\frac{1}{|\mathcal{I}|}\sum_{i\in\mathcal{I}}\text{Prec}_{i}@K$
$\text{DCG}_{u}@K=\sum_{i=1}^{K}\frac{2^{rel_{i,u}}-1}{\log_{2}(i+1)}$ $\text{DCG}_{i}@K=\sum_{j=1}^{K}\frac{2^{rel_{j,i}}-1}{\log_{2}(j+1)}$
$\text{IDCG}_{u}@K=\sum_{i=1}^{K}\frac{1}{\log_{2}(i+1)}$ $\text{IDCG}_{i}@K=\sum_{j=1}^{K}\frac{1}{\log_{2}(j+1)}$
$\text{NDCG}_{u}@K=\frac{\text{DCG}_{u}@K}{\text{IDCG}_{u}@K}$ $\text{NDCG}_{i}@K=\frac{\text{DCG}_{i}@K}{\text{IDCG}_{i}@K}$
$\text{NDCG}_{\text{user}}@K=\frac{1}{|\mathcal{U}|}\sum_{u\in\mathcal{U}}\text{NDCG}_{u}@K$ $\text{NDCG}_{\text{item}}@K=\frac{1}{|\mathcal{I}|}\sum_{i\in\mathcal{I}}\text{NDCG}_{i}@K$
$\text{AP}_{u}@N=\sum_{k=1}^{N}P(k)\cdot\text{rel}_{u}(k)/\min(m,N)$ $\text{AP}_{i}@N=\sum_{k=1}^{N}P(k)\cdot\text{rel}_{i}(k)/\min(m,N)$
$\text{MAP}_{\text{user}}@N=\frac{1}{|\mathcal{U}|}\sum_{u\in\mathcal{U}}\text{AP}_{u}@N$ $\text{MAP}_{\text{item}}@N=\frac{1}{|\mathcal{I}|}\sum_{i\in\mathcal{I}}\text{AP}_{i}@N$
@@ 4.2 Results
User-centric metrics measure the relevance from user’s perspective i.e., whether the relevant items (the user likes or interacts with) are present in the top $K$ recommendations. Users typically expect to see a small number of recommendations (e.g., $5$ or $10$), hence we take $K=5,10,20,50$ to assess how well the system can provide personalized and focused recommendations. Larger values of $K$, such as $20$ or $50$, are useful for evaluating the system’s ability to cover a wider range of user preferences or measure performance in suggesting diverse options and catering to different user tastes. These metrics primarily focus on user satisfaction and do not capture the effectiveness of forgetting specific items. Hence, for item-unlearning, we introduce several item-centric metrics as well.
Item-centric metrics measure the relevance of the recommended items themselves i.e., how well the recommended items meet unlearning criteria, regardless of which users receive them. We take $K=500,1000,1500$ to ensure that all items in the forget set $\mathcal{D}_{f}$ are considered for evaluating performance. We propose them in Table 6.
User Unlearning. Table 2 presents the results of unlearning $5\%$ users. The Gold model, which has never seen the forget set, falls in performance on this set as expected. However, it retains good performance on the test and validation sets. On the validation set, MMRecUn consistently achieves better recall scores, outperforming AmUn by $29.6\%$, $49.85\%$, and $34.75\%$ on the three datasets (baby, sports, and clothing, respectively), showcasing best unlearning generalizability. Furthermore, on the test set, MMRecUn maintains closest recall scores, surpassing AmUn by $26.69\%$, $46.93\%$, and $36.13\%$, demonstrating best retention fidelity. On the forget set, MMRecUn achieves better recall scores, outperforming AmUn by $3.8\%$ and $43.75\%$ on two datasets and slightly falling back by $2.08\%$ on the clothing dataset, indicating high unlearning specificity. Similar trends follow for other metrics.
Item Unlearning Table 3 presents the results of unlearning $5\%$ items. On the validation set, the recall scores differ by $2.21\%$, $0.65\%$, and $0.41\%$ on the three datasets, showcasing comparable unlearning generalizability. Similarly, on the test set, the recall scores differ by $1.43\%$, $0.13\%$, and $0.10\%$, demonstrating comparable retention fidelity. However, MMRecUn achieves good recall scores, outperforming AmUn by $10.83\%$ and $8.45\%$ on two datasets and slightly falling back by $3.43\%$ on the sports dataset, indicating high unlearning specificity.
@@
Tuning hyper-parameter $\mathbf{\alpha}$. We evaluate the impact of $\alpha$ on the balance between adaptation and reform in the unlearning process, experimenting with $\alpha$ values from $0.001$ to $0.9$ (Table 5). Lower $\alpha$ values ($0.1$-$0.2$) effectively adapt to unlearning the forget set with minimal impact on retain set recall, and require fewer epochs for convergence, indicating an optimal balance. As $\alpha$ increases ($0.3$-$0.9$), the model emphasizes more on the reform loss, needing more epochs to forget and adversely affecting retain set recall. Extremely low $\alpha$ values ($<0.01$) hinder effective forgetting, also increasing epochs needed and negatively impacting retain set. The optimal $\alpha$ depends on the specific task’s forgetting severity and time constraints: higher $\alpha$ may ensure stronger forgetting with minor recall loss, while lower $\alpha$ enables faster convergence.
Additional results, including NDCG and MAP metrics, as well as stress testing analysis, are detailed in Appendix A.7.

### v3

@@ Abstract
This paper introduces MMRecUn, the first approach known to us for unlearning in MMRS and unlearning item data. Given a trained RS model, MMRecUn employs a novel Reverse Bayesian Personalized Ranking (BPR) objective to enable the model to forget marked data. The reverse BPR attenuates the impact of user-item interactions within the forget set, while the forward BPR reinforces the significance of user-item interactions within the retain set. Our experiments demonstrate that MMRecUn outperforms baseline methods across various unlearning requests when evaluated on benchmark MMRS datasets. MMRecUn achieves recall performance improvements of up to 49.85% compared to baseline methods and is up to 1.3× faster than the Gold model, which is trained on retain set from scratch. MMRecUn offers significant advantages, including superiority in removing target interactions, preserving retained interactions, and zero overhead costs compared to previous methods.
Codehttps://github.com/MachineUnlearn/MMRecUN
Extended versionhttps://arxiv.org/abs/2405.15328
@@ 1 Introduction
Background and Related Work. Given these benefits, one might ask: Can unlearning methods for uni-modal systems be adapted for the unique challenges of multi-modal systems? Previous unlearning methods (Nguyen et al. 2022) face the following challenges: ❶ MMRS integrate diverse user-item data, including images, text, and behavior, into a unified convolutional graph. This is incompatible with the matrices, latent factors, feature representations, and temporal sequences used in matrix factorization (Liu et al. 2023a; Xu et al. 2022; Liu et al. 2022; Zhang et al. 2023a), collaborative filtering (Li et al. 2024; Schelter, Ariannezhad, and de Rijke 2023), and sequential RS (Ye and Lu 2023), respectively. ❷ For methods that do use graph-based representations, integrating diverse modalities is challenging. Because structure and feature embeddings are tightly integrated, unlearning in one modality has cascading effects on the other (Cheng and Amiri 2023). ❸ Approximate unlearning methods (You et al. 2024; Zhang et al. 2023b; Li et al. 2023b) involve expensive operations, such as inverting Hessian matrix or the Fisher Information Matrix. These computations are costly and impractical for MMRS having a large number of feature dimensions. ❹ Exact unlearning methods (Chen et al. 2022a; Li et al. 2023a) adapt SISA (Bourtoule et al. 2021) to split the dataset into shards. This disrupts the graph structure and degrades performance within each shard due to limited data and poor data heterogeneity (Koch and Soll 2023) by $10-30\%$ (You et al. 2024). The aggregation step, to preserve data structure and aggregate performance across shards, introduces significant overhead costs, impacting both training and inference, increasing proportionally with the number of shards (Ramezani et al. 2021). In worst case, when data that needs to be forgotten come from multiple shards, efficiency falls to the level of retraining from scratch. ❺ While unlearning might appear efficient by retraining only specific shards, the process of setting up and aggregating these shards introduces additional, unaccounted overhead. If the data is initially partitioned by user dimensions, it cannot be re-split for item unlearning. Consequently, simultaneous unlearning of both user and item data becomes impossible.
These constraints underscore the need for methods specifically designed for unlearning in MMRS and for unlearning item data. The challenges and previous works, including  (Cheng and Amiri 2023; Yuan et al. 2023; Li et al. 2023c; Chen et al. 2024; Ganhör et al. 2022; Xin et al. 2024; Wang et al. 2024; Sinha, Mandal, and Kankanhalli 2023; Tarun et al. 2023b; Chundawat et al. 2023b; Tarun et al. 2023a; Chundawat et al. 2023a), provide valuable insights into this domain.
@@ 3 Proposed MMRecUn Method
Traditionally, there are two stages for $M(\cdot,\varphi)$. In training, $M$ encodes collaborative signals inherent in the user-item interaction matrix $\mathcal{Y}$. The optimization process minimizes the BPR loss (Rendle et al. 2012):
@@
Time Complexity for unlearning LightGCN with BPR and auxiliary contrastive loss using MMRecUn is $T=\mathcal{O}\left(\mathcal{M}\cdot d\cdot|\mathcal{E}|\cdot d_{m}+|\mathcal{U}|\cdot|\mathcal{I}|\cdot K+d^{2}+|\mathcal{M}|\cdot d\right)$, where $\mathcal{M}$ is the number of modalities, $d$ is the embedding dimension, $|\mathcal{E}|$ is the number of edges, $d_{m}$ is the feature dimension for modality $m$, $|\mathcal{U}|$ is the number of users, $|\mathcal{I}|$ is the number of items, and $K$ is the number of negative samples per user-item pair.
@@ 4.1 Experimental Setup
Datasets and baselines. We select Mgcn (Yu et al. 2023) as our base model for several reasons: ❶ It is the state-of-the-art MMRS with BPR. ❷ Mgcn represents BPR-based models well, making improvements likely to generalize. ❸ Unlike traditional methods like Collaborative Filtering and Matrix Factorization, it handles multi-modal features through graph convolutional representations. ❹ Other loss functions for uni-modal systems don’t address multi-modal complexities, and those that do, perform weaker than Mgcn. So, focusing on Mgcn ensures our approach is relevant.
We use three distinct categories within the Amazon dataset (Hou et al. 2024) that are used to benchmark MMRS. The model trained on the retain set (named as the Gold model) from scratch is used as baseline. We prepare another baseline by re-purposing the image classification unlearning AmUn (Graves, Nagisetty, and Ganesh 2021) for MMRS.
@@
Evaluation metrics To evaluate a given model $M(\cdot,\varphi)$, we need ❶ model predictions: ranked list of user-item pairs; ❷ the ground truth: user-item interactions $\mathcal{Y}$; and, ❸ $K$: number of the top recommendations to consider. To capture a variety of aspects of performance, we use ❶ predictive metrics, that reflect the “correctness”, how well $M(\cdot,\varphi)$ finds relevant items such as Recall at K and Precision at K; and ❷ ranking metrics, that reflect ranking quality, how well $M(\cdot,\varphi)$ can sort items from more relevant to less relevant. NDCG considers both relevance and position of items in ranked list. MAP measures the average Precision across different Recall levels for a ranked list.
Settings and Hyper-parameters All experiments are performed on $4$x NVIDIA RTX2080 ($32$GB). We use the identical settings as in the case of Mgcn for training the original model $M(\cdot,\varphi)$. The data interaction history of each user is split $8:1:1$ for training, validation and testing. The training set is used to create retain set and forget set. We initialize the embedding with Xavier initialization of dimension $64$, set the regularization coefficient $\lambda_{\varphi}=10^{-4}$, and batch size $B=2048$. For the self-supervised task, we set the temperature to $0.2$. For convergence, the early stopping and total epochs are fixed at $20$ and $1000$, respectively. We use Recall@20 on the validation data as the training-stopping indicator following (Zhang et al. 2022). We conducted experiments five times using different seed values and report the standard deviation values for item unlearning in Table 3. For the other types of unlearning, the standard deviation values were found to be negligible and are therefore omitted.
@@
Metric User-Centric Item-Centric
$\text{Recall}@K$ $\frac{|\text{Relevant}_{u}\cap\text{Retrieved}_{u}|}{|\text{Relevant}_{u}|}$ $\frac{\sum_{j=1}^{K}\text{rel}_{j,i}}{|\text{Users}_{i}|}$
$\text{Recall}@K$ $\frac{1}{|\mathcal{U}|}\sum_{u\in\mathcal{U}}\text{Recall}_{u}@K$ $\frac{1}{|\mathcal{I}|}\sum_{i\in\mathcal{I}}\text{Recall}_{i}@K$
$\text{Prec}@K$ $\frac{|Rel_{u}\cap Rec_{u}|}{|Rec_{u}|}$ $\text{Prec}_{i}@K=\frac{\sum_{j=1}^{K}\text{rel}_{j,i}}{K}$
$\text{Precision}@K$ $\frac{1}{|\mathcal{U}|}\sum_{u\in\mathcal{U}}\text{Prec}_{u}@K$ $\frac{1}{|\mathcal{I}|}\sum_{i\in\mathcal{I}}\text{Prec}_{i}@K$
$\text{DCG}@K$ $\sum_{i=1}^{K}\frac{2^{rel_{i,u}}-1}{\log_{2}(i+1)}$ $\sum_{j=1}^{K}\frac{2^{rel_{j,i}}-1}{\log_{2}(j+1)}$
$\text{IDCG}@K$ $\sum_{i=1}^{K}\frac{1}{\log_{2}(i+1)}$ $\sum_{j=1}^{K}\frac{1}{\log_{2}(j+1)}$
$\text{NDCG}@K$ $\frac{\text{DCG}_{u}@K}{\text{IDCG}_{u}@K}$ $\frac{\text{DCG}_{i}@K}{\text{IDCG}_{i}@K}$
$\text{NDCG}@K$ $\frac{1}{|\mathcal{U}|}\sum_{u\in\mathcal{U}}\text{NDCG}_{u}@K$ $\frac{1}{|\mathcal{I}|}\sum_{i\in\mathcal{I}}\text{NDCG}_{i}@K$
$\text{AP}@N$ $\sum_{k=1}^{N}\frac{P(k)\cdot\text{rel}_{u}(k)}{\min(m,N)}$ $\sum_{k=1}^{N}\frac{P(k)\cdot\text{rel}_{i}(k)}{\min(m,N)}$
$\text{MAP}@N$ $\frac{1}{|\mathcal{U}|}\sum_{u\in\mathcal{U}}\text{AP}_{u}@N$ $\frac{1}{|\mathcal{I}|}\sum_{i\in\mathcal{I}}\text{AP}_{i}@N$
@@ 4.2 Results
User-centric metrics measure the relevance from user’s perspective i.e., whether the relevant items (the user likes or interacts with) are present in the top $K$ recommendations. Users typically expect to see a small number of recommendations (e.g., $5$ or $10$), hence we take $K=5,10,20,50$ to assess how well the system can provide personalized and focused recommendations. Larger values of $K$, such as $20$ or $50$, are useful for evaluating the system’s ability to cover a wider range of user preferences or measure performance in suggesting diverse options and catering to different user tastes. These metrics primarily focus on user satisfaction and do not capture the effectiveness of forgetting specific items. Hence, we introduce several item-centric metrics.
Item-centric metrics (Table 6) measure the relevance of the recommended items themselves i.e., how well the recommended items meet unlearning criteria, regardless of which users receive them. We take $K=500,1000,1500$ to ensure that all items in the forget set $\mathcal{D}_{f}$ are considered for evaluating performance.
User Unlearning. Table 2 shows the results of unlearning $5\%$ users. The Gold model, which has never seen the forget set, falls in performance on this set as expected. However, it retains good performance on the test and validation sets. On the validation set, MMRecUn consistently achieves better recall scores, outperforming AmUn by $29.6\%$, $49.85\%$, and $34.75\%$ on the three datasets (baby, sports, and clothing, respectively), showcasing best unlearning generalizability. On the test set, MMRecUn maintains closest recall scores, surpassing AmUn by $26.69\%$, $46.93\%$, and $36.13\%$, demonstrating best retention fidelity. On the forget set, MMRecUn achieves better recall scores, outperforming AmUn by $3.8\%$ and $43.75\%$ on two datasets and slightly falling back by $2.08\%$ on the clothing dataset, indicating high unlearning specificity. Similar trends follow for other metrics.
Item Unlearning Table 3 shows the results of unlearning $5\%$ items. On the validation set, the recall scores differ by $2.21\%$, $0.65\%$, and $0.41\%$ on the three datasets, showcasing comparable unlearning generalizability. Similarly, on the test set, the recall scores differ by $1.43\%$, $0.13\%$, and $0.10\%$, demonstrating comparable retention fidelity. However, MMRecUn achieves good recall scores, outperforming AmUn by $10.83\%$ and $8.45\%$ on two datasets and slightly falling back by $3.43\%$ on the sports dataset, indicating high unlearning specificity.
@@
Tuning hyper-parameter $\mathbf{\alpha}$. We assess how $\alpha$ balances adaptation and reform in unlearning, testing values from $0.001$ to $0.9$ (Table 5). Lower $\alpha$ values ($0.1$-$0.2$) effectively unlearn the forget set with minimal impact on retain set recall, and require fewer epochs for convergence, indicating an optimal balance. As $\alpha$ increases ($0.3$-$0.9$), the model emphasizes more on the reform loss, needing more epochs to forget, adversely affecting retain set recall. Extremely low $\alpha$ values ($<0.01$) hinder effective forgetting, also increasing epochs and negatively impacting retain set. The optimal $\alpha$ depends on the specific task’s forgetting severity and time constraints: higher $\alpha$ may ensure stronger forgetting with minor recall loss, while lower $\alpha$ enables faster convergence.